OpenStudy (sarahhanish):

What is the slope of the line whose equation is 12=4x−6y FOR THE LOVE OF DONALD TRUMP ? Enter your answer in the boxes.

OpenStudy (sarahhanish):

may you help?

OpenStudy (steve816):

Heyyy, do you think you can get the equation to y = mx+b form?

OpenStudy (sarahhanish):

oh....hey steve...

OpenStudy (sarahhanish):

mmm, ill try

OpenStudy (sarahhanish):

y=6x4? ughhh idk

OpenStudy (sarahhanish):

ugh im left alone with my stupidness

OpenStudy (steve816):

It's ok, you can do it! 

zepdrix (zepdrix):

@marcelie lolol

OpenStudy (steve816):

12 = 4x-6y -6y = -4x + 12 \[\huge y=\frac{ 2 }{ 3 }x-2\]

OpenStudy (sarahhanish):

how did you even get that?!?!?! UGH WIHS MATH WOULD JUST DIEE

OpenStudy (steve816):

We have the equation in y=mx+b form! yayy

OpenStudy (marcelie):

omg hahahaha Ariel!! :3

OpenStudy (steve816):

I know I know

OpenStudy (sarahhanish):

BUT-BUT UGHHHH i dont get it how you did it!

OpenStudy (marcelie):

sarah just for curiosity was ur name misty? loool

OpenStudy (sarahhanish):

misty?

OpenStudy (steve816):

huh, you know mistyfox?

OpenStudy (sarahhanish):

nope

OpenStudy (steve816):

Oh, nevermind then.

OpenStudy (sarahhanish):

why?

OpenStudy (steve816):

nothing

OpenStudy (steve816):

Anyways back to math!

OpenStudy (steve816):

The slope would be 2/3

OpenStudy (sarahhanish):

but how u know?

OpenStudy (steve816):

Because we got the equation into y = mx+b form, and m is the slope :)